OpenStudy (anonymous):

16p = 208 p =----- a0

OpenStudy (anonymous):

please help me

OpenStudy (lgbasallote):

divide 208 by 16..what do you get?

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

144 = m - 98 m = a0

OpenStudy (lgbasallote):

add 98 to both sides

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

= 300 n = a0

OpenStudy (anonymous):

pleaseheeeelp meeee

OpenStudy (anonymous):

n/12= 300 n = a0